OpenStudy (anonymous):

find the area of a circle whose circumference is 8 pie m? a. 4 pie m squared b. 8 pie m squared c. 16 pie m squared d. 64 pie m squared

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

@AJ01

OpenStudy (rational):

start by finding the radius of circle using given info

OpenStudy (rational):

remember circumference of circle formula ?

OpenStudy (anonymous):

not rly

OpenStudy (rational):

must be there in ur notes, check once

OpenStudy (anonymous):

dont have any

OpenStudy (anonymous):

so its 2 times 3.14 times 8 right?

OpenStudy (rational):

leave pie as pie, dont replace it with 3.14

OpenStudy (rational):

\[2\pi r = 8\pi\] solve \(r\) first

OpenStudy (anonymous):

24?

OpenStudy (rational):

\[2\pi r = 8\pi\] cancel \(2\pi\) both sides \[r = 4\]

OpenStudy (rational):

plug that in area of circle formula

OpenStudy (anonymous):

32

OpenStudy (anonymous):

use the same equations area= pi*r^2, circ= 2*pi* r you have the circ so find r using 2nd equation then use the value of r in the second equation....

OpenStudy (anonymous):

circ= 2*pi*r= 8 pi => r=(8 pi)/(2 pi)=4 area= pi r^2= pi (4)^2= 16 pi

OpenStudy (anonymous):

so is that the final answer???

OpenStudy (anonymous):

C

OpenStudy (anonymous):

yes